# NFL-BA: Near-Field Light Bundle Adjustment for SLAM in Dynamic Lighting

### v3

@@ 5.1 Evaluation Setting
C3VD. The C3VD dataset c3vd (CC BY-NC-SA 4.0) was created using a phantom colon with synthetic materials to simulate realistic tissue geometry. The endoscopy video was captured by a surgeon who performs different endoscopy procedures on the phantom colon with a real endoscope capturing RGB images coupled with corresponding depth maps. We focus on 8 sequences ranging from 70 to 800 frames from different regions of the colon, for more details, please see supplementary. We evalute using both ground truth and predicted depths.
Colon10K. To test generalization in real-world clinical endoscopy settings, we evaluate on Colon10K Colon10k , a large-scale video dataset without depth or pose supervision. Videos are sampled from actual
Figure 5: Camera tracking improvement over EndoGSLAM Wan_EndoGSLAM_MICCAI2024 . Replacing the Photo-BA loss (in blue) with NFL-BA loss (in red) significantly improves camera tracking for different depth initialization. Average tracking error ATEt for each sequence is reported in the inset. (zoom for details)
@@
Self-Captured Indoor Scenes. To study the role of near-field lighting in a controlled non-clinical setting, we capture a dataset of four indoor scenes (Guitars, Porch, Pool, and Stairs) using an RGB-D camera. Scenes include objects with varied geometry and reflectance (diffuse, specular), imaged under dynamic motion. Scenes are captured using a co-located point light source mounted to camera. Ground-truth camera trajectories were recorded via motion capture, but no reference point clouds are available; hence, we report only trajectory error (ATEt) and perceptual quality (LPIPS), omitting Chamfer distance. Details and additional visualizations are included in the supplementary.
Metrics. For evaluation, we basically followed other neural rendering SLAM algorithms MonoGS ; Wan_EndoGSLAM_MICCAI2024 . For tracking performance, we measure the root mean square error of the Absolute Trajectory Error (ATE) for both translation and rotation across all frames. Translation error $\text{ATE}_{t}$ is in millimeters (mm) for the endoscopy scenes and meters (m) for the in-door scenes. and rotation error $\text{ATE}_{r}$ is in degrees. To assess the mapping quality, we use the Chamfer distance from ground truth point clouds to the nearest points in the estimated point clouds NormDepth , for more details please see supplementary. In addition, we evaluate rendering quality using the Learned Perceptual Image Patch Similarity (LPIPS) LPIPS . We note that for many endoscopic SLAM applications, tracking and mapping accuracies are more important than photorealism of the rendered images, unlike many indoor or outdoor scenes.
Computational costs We trained all models on a single NVIDIA RTX A6000 GPU. The per-scene optimization takes $\sim$1 FPS. For more information on runtime speed, please see supplemental.
@@ 5.2 Evaluation on C3VD Endoscopy Data
Table 2: Quantitative Evaluation on the C3VD c3vd dataset with depth maps estimated by SOTA techniques, PPSNet PPSNet and DA-Hybrid. Replacing Photometric BA with NFL-BA significantly improves tracking for both MonoGS MonoGS and EndoGS EndoGS , and mapping and rendering quality for MonoGS MonoGS . Note that SOTA performance for each of the tracking, mapping, and rendering metrics is observed when NFL-BA is used.
[…]
Method BA PPS-Net DaHybrid PPS-Net DaHybrid PPS-Net DaHybrid PPS-Net DaHybrid
EndoGSLAM, MICCAI’24 Photo 3.03 6.67 1.73 2.26 1.23 2.12 0.39 0.43
NFL 2.62 3.91 1.24 1.58 1.25 2.39 0.39 0.42
MonoGS, CVPR’24 Photo 3.48 4.63 1.70 1.69 1.59 1.34 0.56 0.52
NFL 2.18 2.35 1.65 1.14 0.99 1.13 0.53 0.52
Figure 6: (a) Camera tracking improvement over MonoGS MonoGS . Replacing the Photo BA loss (in blue) with NFL-BA loss (in red) significantly improves camera tracking for different depth initialization. Average tracking error ATEt for each sequence is reported in the inset. (b) Reconstructed point clouds using MonoGS MonoGS show that NFL-BA improves coverage and density while reducing scatter compared to Photometric BA, as measured by Chamfer distance. (zoom for details)
@@
SLAM with predicted depth map. Under realistic conditions with estimated depths, NFL-BA’s impact is even more pronounced. In Tab. LABEL:tab:pred_depth we replace Photometric BA loss with NFL-BA loss for MonoGS MonoGS and EndoGSLAM Wan_EndoGSLAM_MICCAI2024 for depth maps we use PPSNet PPSNet , a state-of-the-art monocular depth estimation algorithm for endoscopy, and fine-turned general-purpose depth estimator, which we will call it as DA-Hybrid - DepthAnythingdepthanything with DINOv2 encoder dinov2 . NFL-BA significantly improves camera localization ($ATE_{t}$) and camera rotation ($ATE_{r}$) for both MonoGS MonoGS and EndoGSLAM Wan_EndoGSLAM_MICCAI2024 while producing similar rendering quality. For example, camera localization for MonoGS is improved by 37% for PPSNet and 49% for DA-Hybrid depth initialization. Mapping accuracy of MonoGS also improves by 37% for PPSNet and 16% for DA-Hybrid depth maps. Overall, these results demonstrate that NFL-BA can compensate for noisy depth estimation and improves performance. Across all four metrics, for tracking, mapping, and rendering, the SOTA performance on the C3VD dataset is in fact achieved when NFL-BA loss is used in the SLAM framework.

### v4

@@ 5.1 Evaluation Setting
Table 2: Quantitative evaluation on the C3VD Bobrow et al. (2023) dataset using depth maps estimated by SOTA techniques, PPSNet Paruchuri et al. (2024) and DA-Hybrid Yang et al. (2024a); Oquab et al. (2024). Replacing Photometric BA with NFL-BA significantly improves tracking for both MonoGS Matsuki et al. (2024) and EndoGSLAM Wang et al. (2024), and mapping and rendering quality for MonoGS Matsuki et al. (2024). Note that SOTA performance for each of the tracking, mapping, and rendering metrics is observed when NFL-BA is used.
Tracking Mapping Rendering
[…]
PPS-Net Photo 3.03 1.73 1.23 0.39
EndoGSLAM Wang et al. (2024) PPS-Net NFL 2.62 1.24 1.25 0.39
MICCAI’24 DA-Hybrid Photo 6.67 2.26 2.12 0.43
DA-Hybrid NFL 3.91 1.58 2.39 0.42
PPS-Net Photo 3.48 1.70 1.59 0.56
MonoGS Matsuki et al. (2024) PPS-Net NFL 2.18 1.65 0.99 0.53
CVPR’24 DA-Hybrid Photo 4.63 1.69 1.34 0.52
DA-Hybrid NFL 2.35 1.14 1.13 0.52
@@
C3VD. The C3VD dataset Bobrow et al. (2023) (CC BY-NC-SA 4.0) was created using a phantom colon with synthetic materials to simulate realistic tissue geometry. Colon10K. To test generalization in real-world clinical endoscopy settings, we evaluate on Colon10K Ma et al. (2021), a large-scale video dataset without depth or pose supervision. Videos are sampled from actual The endoscopy video was captured by a surgeon who performs different endoscopy procedures on the phantom colon with a real endoscope capturing RGB images coupled with corresponding depth maps. We focus on 8 sequences ranging from 70 to 800 frames from different regions of the colon, for more details, please see supplementary. We evalute using both ground truth and predicted depths.
Figure 5: Camera tracking improvement over EndoGSLAM Wang et al. (2024). Replacing the Photo-BA loss (in blue) with NFL-BA loss (in red) significantly improves camera tracking for different depth initialization. Average tracking error ATEt for each sequence is reported in the inset. (zoom for details)
@@
Self-Captured Indoor Scenes. To study the role of near-field lighting in a controlled non-clinical setting, we capture a dataset of four indoor scenes (Guitars, Porch, Pool, and Stairs) using an iPhone 15 Pro. Scenes include objects with varied geometry and reflectance (diffuse, specular), imaged under dynamic motion. Scenes are captured using a co-located point light source mounted to camera. Ground-truth camera trajectories were recorded via motion capture, but no reference point clouds are available; hence, we report only trajectory error (ATEt) and perceptual quality (LPIPS), omitting Chamfer distance. Details and additional visualizations are included in the supplementary.
Metrics. For evaluation, we basically followed other neural rendering SLAM algorithms Matsuki et al. (2024); Wang et al. (2024). For tracking performance, we measure the root mean square error of the Absolute Trajectory Error (ATE) for both translation and rotation across all frames. Translation error $\text{ATE}_{t}$ is in millimeters (mm) for the endoscopy scenes and meters (m) for the in-door scenes. And rotation error $\text{ATE}_{r}$ is in degrees. To assess the mapping quality, we use the Chamfer distance from ground truth point clouds to the nearest points in the estimated point clouds Wang et al. (2023b), for more details please see supplementary. In addition, we evaluate rendering quality using the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018). We note that for many endoscopic SLAM applications, tracking and mapping accuracies are more important than photorealism of the rendered images, unlike many indoor or outdoor scenes.
Computational costs We trained all models on a single NVIDIA RTX A6000 GPU. The per-scene optimization takes $\sim$1 FPS. For more information on runtime speed, please see supplementary materials.
Figure 6: Camera tracking improvement over MonoGS Matsuki et al. (2024). Replacing the Photo BA loss (in blue) with NFL-BA loss (in red) significantly improves camera tracking for different depth initialization. Average tracking error ATEt for each sequence is reported in the inset.
Figure 7: Reconstructed point clouds using MonoGS Matsuki et al. (2024) show that NFL-BA improves coverage and density while reducing scatter compared to Photometric BA, as measured by Chamfer distance.
@@ 5.2 Evaluation on C3VD Endoscopy Data
SLAM with predicted depth map. Under realistic conditions with estimated depths, NFL-BA’s impact is even more pronounced. In Tab. 2 we replace Photometric BA loss with NFL-BA loss for MonoGS Matsuki et al. (2024) and EndoGSLAM Wang et al. (2024) for depth maps we use PPSNet Paruchuri et al. (2024), a state-of-the-art monocular depth estimation algorithm for endoscopy, and fine-turned general-purpose depth estimator, which we will call it as DA-Hybrid - DepthAnythingYang et al. (2024a) with DINOv2 encoder Oquab et al. (2024). NFL-BA significantly improves camera localization ($ATE_{t}$) and camera rotation ($ATE_{r}$) for both MonoGS Matsuki et al. (2024) and EndoGSLAM Wang et al. (2024) while producing similar rendering quality. For example, camera localization for MonoGS is improved by 37% for PPSNet and 49% for DA-Hybrid depth initialization. Mapping accuracy of MonoGS also improves by 37% for PPSNet and 16% for DA-Hybrid depth maps. Overall, these results demonstrate that NFL-BA can compensate for noisy depth estimation and improves performance. Across all four metrics, for tracking, mapping, and rendering, the SOTA performance on the C3VD dataset is in fact achieved when NFL-BA loss is used in the SLAM framework.
@@ 6 Conclusions
7 Acknowledgments
This work is supported by a National Institute of Health (NIH) project #R21EB035832 "Next-gen 3D Modeling of Endoscopy Videos" and #R21EB037440 “Gen-AI Airway Simulator for 3D Endoscopy”. We also thank Stephen M. Pizer, Ron Alterovitz, and Dr. Sarah McGill for helpful discussions during the project.
Supplementary Materials
Overview of Appendices
Our appendices contain the following additional details:
  • Appendix A: Dataset Processing and Collection
  • Appendix B: Implementation Details and Computational Costs
  • Appendix C: Point Clouds and Metrics
  • Appendix D: Long Sequence Validation
  • Appendix E: Evaluation of Center-Crop Baseline
  • Appendix F: Lighting model and Angular Attenuation
  • Appendix G: Results on Endoscopy Datasets
Appendix A Dataset Processing and Collection
C3VD. We test our method on a subset of 8 videos with at least one video from each section of the colon, with varying camera motion, and anomalies. We choose these 8 videos from the test split of PPSNet to avoid any bias when the SLAM is initialized with PPSNet predicted depth map. The names of the sequences are as follows: cecum_t1_a, cecum_t2_a, cecum_t3_a, sigmoid_t3_a, desc_t4_a_p2, trans_t2_a, trans_t3_a, and trans_t4_a.
All images were cropped to remove any artifacts resulting from fish-eye correction then downscale and crop the images. Specifically, we resize each image to have a height of 384 pixels while maintaining the aspect ratio, then crop the central region to obtain a 384×384 pixel image.
Colon10k. Since Colon10K provides no ground‐truth depths, we compute per-frame estimates with PPSNet. Each image is center-cropped and uniformly resized to 384 × 384 px to match our SLAM input requirements.
Indoor Self Captures. We recorded four indoor scenes using an iPhone 15 with LiDAR, capturing synchronized RGB (1440 × 1920 px) and depth (256 × 192 px) streams. Raw RGB frames are downscaled to 256 × 192 px to align with the depth map resolution. Depth maps are stored as 16-bit values up to 10 m. We logged camera poses via Apple’s ARKit framework, code for our custom capture app and preprocessing scripts will be released alongside the dataset.
Appendix B Implementation Details and Computational Costs
As mentioned in the main paper, we ran all models on a single NVIDIA RTX A6000 GPU. The per-scene optimization takes approximately 1 FPS. We found that NFL-BA only reduces the fps runtime by a small amount on the C3VD dataset.
Table 4: Runtime: We evaluate the frames per second (FPS) for all methods on the C3VD dataset.
Method Depth Photo-BA NFL-BA
NICE-SLAM Oracle $\ll$ 1 $\ll$ 1
PPSNet $\ll$ 1 $\ll$ 1
EndoGSLAM Oracle 1.79 1.35
PPSNet 1.53 1.22
DPT-Hybrid 1.20 0.90
MonoGS Oracle 1.38 1.09
PPSNet 1.06 0.93
DPT-Hybrid 0.99 0.83
NICE-SLAM. Since the scene is encoded using neural networks, we extract normals from the occupancy grid, as described in Sec. 4.2 to calculate the shading term. NICE-SLAM requires a well-defined bounding box which we obtained from the ground truth point clouds (see appendix C). We also used the default loss weights of NICE-SLAM, setting $\lambda_{ren}$ to 0.5 during tracking and 0.2 during mapping, and $\lambda_{geo}$ to 1 in both phases.
EndoGSLAM. The main difference is the weight map $M_{t}$ in the bundle adjustment loss (Eq. 3) to exclude over-exposed pixels that can arise in endoscopy-specific lighting conditions. Furthermore, given that the shading term $PPS(\cdot)$ is sensitive to depth scales, we rescaled the depth maps so that their maximum values are approximately 5. Notably, scaling the depth maps did not improve baseline performance (when using PPS depth maps, the average $ATE_{t}$ went from 3.03 to 3.38). We used the default loss weights of EndoGSLAM, $\lambda_{ren}$; $\lambda_{geo}$, set to 0.5 and 1 during tracking and 1 and 1 during mapping, respectively.
MonoGS. To integrate our method, we treat the Gaussian color features as albedo features and multiply them with the shading term before rasterization, and then we use the rendered output colors for bundle adjustment. For all input depths, we set $\lambda_{ren}$ and $\lambda_{geo}$ to 0.8 and 0.5, respectively.
Appendix C Point Clouds and Metrics
Chamfer Distances. Since ground truth point clouds are unavailable for C3VD, we generate them by unprojecting 2D images into 3D space with the correct camera configuration and the oracle depth maps, provided in the C3VD dataset. For neural fields-based SLAM, we use the vertices of the output meshes as the estimated point clouds, while for Gaussian Splatting-based SLAMs, we use the Gaussian positions. For point cloud alignment, we use Coherent Point Drift Myronenko and Song (2010) and the Chamfer distances are also in millimeters.
Coloring Point Clouds. We use extracted point clouds from 3D Gaussian positions for visualization. For colors, we directly use Gaussian color features.
Appendix D Long Sequence Validation
To evaluate NFL-BA’s performance on longer sequence, we test our method on a longer screening video for C3VD, specifically c0_full_t2_v2, which spans over 4,000 frames with ground-truth poses, allowing us to measure cumulative drift. In the main paper, we only evaluate on registered videos with ground truth depth, not screening videos. We show how NFL-BA performs relative to Photo-BA using the MonoGS backbone in Table 5.
NFL-BA matches Photometric BA on short sequences and increasingly outperforms it as trajectory length grows. As a plug-and-play bundle-adjustment loss, NFL-BA enhances long-term robustness under dynamic near-field lighting.
Table 5: ATE_T (cm) on c0_full_t2_v2 at varying lengths. NFL-BA matches or beats Photo-BA on short sequences and shows long-term stability.
BA 500 frames 1,000 frames 2,000 frames 4,000 frames
Photo 2.599 1.937 8.459 14.790
NFL 2.422 2.254 5.742 11.368
Appendix E Evaluation of Center-Crop Baseline
We compared MonoGS Photo-BA optimized on only the central 75% and 50% of each frame against full-frame MonoGS + NFL-BA on two sequences, measuring translational ATE ($\text{ATE}_{T}$), Chamfer distance (CD), and reconstructed point count in Table 6.
Although center-cropping improves camera tracking performance (ATE_T) of MonoGS and gets close to NFL-BA, it results in significantly worse reconstruction in terms of quality (CD) and density (number of points). This is because for camera localization, not all pixels are essential, and focusing only on central pixels can eliminate near-field lighting effects. In contrast, for reconstruction, all pixels matter. This highlights the need for a principled mechanism for handling dynamic near-field lighting, as proposed by NFL-BA.
Table 6: Center-cropping improves trajectory error but reduces map density, while full-frame NFL-BA maintains both accuracy and dense reconstructions.
Method trans_t3_a desc_t4_p2
$\text{ATE}_{T}$ CD Points $\text{ATE}_{T}$ CD Points
MonoGS + NFL-BA 0.26 0.60 28,196 0.13 0.67 8,879
MonoGS (full frame) 0.31 0.76 7,095 0.25 0.76 6,398
MonoGS (75% center) 0.35 1.06 5,761 0.13 0.97 4,329
MonoGS (50% center) 0.40 2.82 1,865 0.15 1.56 2,100
Appendix F Lighting Model and Angular Attenuation
We clarify that NFL-BA models only diffuse reflectance and direct illumination, ignoring specular and subsurface effects. Explicitly modeling these introduces non-differentiable and highly non-convex terms, remaining an open challenge.
To mitigate specular highlights, we apply an intensity mask discarding pixels above 0.9 grayscale intensity. This handles most artifacts but cannot correct reflective surfaces beyond the mask. While effective on matte datasets (C3VD, in-the-wild), degradation occurs on Colon10K with more specular highlights. Future work will address specular reflections under dynamic lighting.
Regarding angular attenuation, $\beta=0$ is justified for tightly collimated endoscopic LEDs, we must validated this for non endoscopy scenes. Using NFL-BA on the MonoGS backbone, we validate various $\beta$ values on two indoor sequences as seen in Table 7
$\beta=0$ provides competitive mean accuracy and low variance. Non-zero values may yield small gains but introduce instability. Given minimal benefit versus tuning cost, we retain $\beta=0$, and plan to explore learning $\beta$ as a per-scene parameter in future work.
Table 7: Ablation of the angular-attenuation coefficient $\beta$ on two indoor sequences, reporting translational ATE and LPIPS.
Scene Metric $\beta$=0.00 $\beta$=0.25 $\beta$=0.50 $\beta$=0.75 $\beta$=1.00
[…]
Appendix G Results on Endoscopy Datasets
For all experiments for each of the slam systems, we report the median of three runs for all tables and figures. We have included per-sequence metrics for the median run below. Additionally, we include point clouds for EndoGSLAM in Figure 10, more can be found on our project page.
Figure 10: EndoGSLAM Point cloud results for 2 sequences: cecum_t1_a_under_review (top) and desc_t4_a_p2_under_review (bottom).
Table 8: Results for sequence cecum_t1_a_under_review
[…]
NICE-SLAM Oracle Photo 2.65 2.82 0.08
NFL 1.39 2.81 0.15
PPS-Net Photo 10.14 2.71 0.51
NFL 4.14 2.75 0.19
EndoGSLAM Oracle Photo 1.39 0.24 0.12
NFL 0.91 0.31 0.16
PPS-Net Photo 2.79 0.60 0.30
NFL 2.93 0.70 0.35
DA-Hybrid Photo 2.64 0.08 0.04
NFL 8.44 2.42 1.21
MonoGS Oracle Photo 1.16 0.39 1.56
NFL 1.22 0.35 0.95
PPS-Net Photo 2.30 0.65 1.19
NFL 2.45 0.76 1.04
DA-Hybrid Photo 5.44 0.30 1.64
NFL 1.09 0.40 1.65
Table 9: Results for sequence cecum_t2_a_under_review
[…]
NICE-SLAM Oracle Photo 8.13 2.19 0.49
NFL 1.15 2.82 0.11
PPS-Net Photo 1.11 2.13 0.11
NFL 7.98 2.82 0.50
EndoGSLAM Oracle Photo 2.32 1.06 0.53
NFL 6.75 2.79 1.40
PPS-Net Photo 4.55 1.18 0.59
NFL 4.55 2.82 1.41
DA-Hybrid Photo 5.66 0.89 0.45
NFL 9.47 2.25 1.13
MonoGS Oracle Photo 4.04 0.34 1.47
NFL 6.84 0.48 1.24
PPS-Net Photo 6.72 2.71 1.74
NFL 6.52 2.73 1.70
DA-Hybrid Photo 5.20 0.66 2.18
NFL 4.26 0.32 1.35
Table 10: Results for sequence cecum_t3_a_under_review
[…]
NICE-SLAM Oracle Photo 3.46 2.82 0.11
NFL 3.42 2.82 0.07
PPS-Net Photo 2.80 2.64 0.17
NFL 3.83 2.59 0.14
EndoGSLAM Oracle Photo 0.75 0.15 0.08
NFL 0.74 0.27 0.14
PPS-Net Photo 0.79 0.16 0.08
NFL 2.18 0.22 0.11
DA-Hybrid Photo 1.45 0.62 0.31
NFL 1.52 0.18 0.09
MonoGS Oracle Photo 0.36 0.16 1.23
NFL 0.87 0.31 0.62
PPS-Net Photo 1.07 0.17 1.12
NFL 1.24 0.14 0.81
DA-Hybrid Photo 1.06 0.33 0.95
NFL 1.16 0.30 0.93
Table 11: Results for sequence desc_t4_a_p2_under_review
[…]
NICE-SLAM Oracle Photo 15.90 2.65 0.59
NFL 8.88 2.75 0.50
PPS-Net Photo 0.87 2.80 0.10
NFL 0.68 2.80 0.10
EndoGSLAM Oracle Photo 0.36 0.56 0.28
NFL 0.26 0.94 0.47
PPS-Net Photo 0.48 0.58 0.29
NFL 1.00 0.83 0.42
DA-Hybrid Photo 0.45 1.16 0.58
NFL 1.80 2.81 1.40
MonoGS Oracle Photo 0.25 1.05 0.76
NFL 0.13 0.98 0.67
PPS-Net Photo 0.49 1.78 0.77
NFL 0.61 1.67 0.70
DA-Hybrid Photo 0.69 2.63 0.80
NFL 0.20 0.97 0.77
Table 12: Results for sequence sigmoid_t3_a_under_review
[…]
NICE-SLAM Oracle Photo 1.04 2.47 0.05
NFL 0.84 2.83 0.05
PPS-Net Photo 9.47 2.79 0.28
NFL 6.60 2.55 0.35
EndoGSLAM Oracle Photo 4.81 1.14 0.57
NFL 4.11 2.69 1.35
PPS-Net Photo 3.13 1.40 0.70
NFL 9.82 2.56 1.29
DA-Hybrid Photo 8.51 2.36 1.18
NFL 8.15 2.80 1.41
MonoGS Oracle Photo 1.44 0.46 1.17
NFL 1.19 2.33 0.67
PPS-Net Photo 1.20 1.53 4.63
NFL 1.22 2.22 0.89
DA-Hybrid Photo 6.88 2.18 1.42
NFL 1.10 0.67 1.33
Table 13: Results for sequence trans_t2_a_under_review
[…]
NICE-SLAM Oracle Photo 0.77 2.83 0.05
NFL 0.94 2.80 0.05
PPS-Net Photo 9.18 2.82 0.26
NFL 9.85 2.81 0.31
EndoGSLAM Oracle Photo 4.21 1.80 0.90
NFL 0.49 2.82 1.41
PPS-Net Photo 10.05 1.66 0.84
NFL 0.90 1.46 0.73
DA-Hybrid Photo 9.14 2.77 1.39
NFL 14.59 2.18 1.10
MonoGS Oracle Photo 14.38 1.42 1.49
NFL 1.51 2.72 0.97
PPS-Net Photo 14.64 1.91 1.52
NFL 3.08 1.02 1.09
DA-Hybrid Photo 15.18 2.12 1.68
NFL 9.51 1.41 1.45
Table 14: Results for sequence trans_t3_a_under_review
[…]
NICE-SLAM Oracle Photo 0.97 2.81 0.15
NFL 6.26 2.79 0.68
PPS-Net Photo 3.78 2.80 0.22
NFL 1.12 2.65 0.13
EndoGSLAM Oracle Photo 0.17 2.70 1.35
NFL 0.20 2.70 1.35
PPS-Net Photo 0.30 2.80 1.40
NFL 0.50 2.81 1.41
DA-Hybrid Photo 0.46 2.81 1.41
NFL 0.33 2.74 1.37
MonoGS Oracle Photo 0.31 2.67 0.76
NFL 0.26 2.66 0.60
PPS-Net Photo 0.33 2.79 0.89
NFL 0.61 2.83 0.82
DA-Hybrid Photo 0.29 2.81 0.91
NFL 0.38 2.71 0.71
Table 15: Results for sequence trans_t4_a_under_review
[…]
NICE-SLAM Oracle Photo 0.39 2.81 0.05
NFL 0.23 2.83 0.05
PPS-Net Photo 7.26 2.77 0.14
NFL 7.88 2.73 0.23
EndoGSLAM Oracle Photo 2.64 1.47 0.74
NFL 1.99 2.02 1.01
PPS-Net Photo 2.02 1.73 0.86
NFL 2.58 1.78 0.89
DA-Hybrid Photo 2.99 2.08 1.04
NFL 10.67 2.72 1.37
MonoGS Oracle Photo 1.23 2.41 0.83
NFL 0.76 2.06 0.62
PPS-Net Photo 1.12 2.10 0.89
NFL 1.70 1.80 0.86
DA-Hybrid Photo 2.30 2.46 1.10
NFL 1.14 2.36 0.85